OpenStudy (anonymous):

a square has a side length of 95. hat is the length of the diagonal of the square? express the answer in simlest form a square has a side length of 95. hat is the length of the diagonal of the square? express the answer in simlest form @Mathematics

OpenStudy (anonymous):

pythagoras: x^2 + x^2 = diagonal^2 2x^2 = diagonal^2 2*(95)^2 = diagonal^2 diagonal = sqrt(2*(95)^2) = 95*sqrt(2)

OpenStudy (anonymous):

 y=sqrt(X^2+x^2)